# ReasonPlan: Unified Scene Prediction and Decision Reasoning for Closed-loop Autonomous Driving

### v1

@@ Abstract
Due to the powerful vision-language reasoning and generalization abilities, multimodal large language models (MLLMs) have garnered significant attention in the field of end-to-end (E2E) autonomous driving. However, their application to closed-loop systems remains underexplored, and current MLLM-based methods have not shown clear superiority to mainstream E2E imitation learning approaches. In this work, we propose ReasonPlan, a novel MLLM fine-tuning framework designed for closed-loop driving through holistic reasoning with a self-supervised Next Scene Prediction task and supervised Decision Chain-of-Thought process. This dual mechanism encourages the model to align visual representations with actionable driving context, while promoting interpretable and causally grounded decision making. We curate a planning-oriented decision reasoning dataset, namely PDR, comprising 210k diverse and high-quality samples. Our method outperforms the mainstream E2E imitation learning method by a large margin of $19\%$ L2 and $16.1$ driving score on Bench2Drive benchmark. Furthermore, ReasonPlan demonstrates strong zero-shot generalization on unseen DOS benchmark, highlighting its adaptability in handling zero-shot corner cases. Code and dataset will be found in https://github.com/Liuxueyi/ReasonPlan.
@@ 1 Introduction
Recent efforts have explored dual-system frameworks [13, 14, 15], LLM distillation for enhancing E2E driving [16, 17], and direct trajectory prediction in textual form [18, 19, 20]. While promising, these approaches predominantly operate in open-loop settings or exhibit suboptimal performance in closed-loop evaluations. This limitation stems from their inability to perform context-aware reasoning and robust planning in closed-loop scenarios, where continuous adaptation to dynamic environments is essential [21]. We conclude three key challenges that limit the full exploitation of MLLMs’ reasoning capabilities: (1) Only unimodal text supervision. Existing methods [22] supervise the perception and decision-making process using only text, resulting in insufficient scene understanding and limited situational awareness. (2) No explicit reasoning process. Prior works [18, 19] utilize multi-turn QAs fine-tuning LLMs to enhance the instruction-following ability, but they fall short of engaging the models’ Chain-of-Thought (CoT) reasoning capabilities. (3) Lack of planning-oriented high-quality reasoning datasets. Consequently, the full potential of MLLMs for E2E closed-loop planning remains underexplored, and current MLLM-based methods have yet to demonstrate clear superiority over imitation learning techniques [1, 2, 23] in closed-loop benchmarks.
@@
From the dataset perspective, we construct a large-scale instruction dataset tailored for closed-loop planning, called PDR, which contains 203,353 training samples and 11,047 testing samples. Using an automated annotation pipeline, PDR captures the complete decision reasoning process in training scenarios on the Bench2Drive [26], including the following stages: Scene Understanding, Traffic Sign Recognition, Critical Object Identification for Risk Assessment, and Meta Action. To ensure the quality and reliability of the dataset, all reasoning steps have undergone thorough human verification. The dataset will be released publicly, serving as a foundation for learning structured and causally grounded decision reasoning.
ReasonPlan achieves a driving score of 64.01 and reduces the L2 error by 16.44% on Bench2Drive, outperforming prior state-of-the-art E2E imitation learning models. Furthermore, it demonstrates strong zero-shot generalization on DOS, highlighting its adaptability in decision-critical corner cases.
Our key contributions are as follows. (1) We propose ReasonPlan, a novel MLLM fine-tuning framework for complex closed-loop driving scenarios. The integration of self-supervised NSP and supervised DeCoT effectively couples the visual and language modalities, enabling a comprehensive decision reasoning process. (2) We construct PDR, a large-scale decision reasoning dataset via an automated annotation pipeline tailored for closed-loop planning, containing 210k diverse and high-quality samples. (3) As shown in Fig. 1, ReasonPlan demonstrates superior performance on the Bench2Drive under both open-loop and closed-loop settings, and shows strong zero-shot generalization on DOS scenarios.
@@ 3.1 Self-supervised Next Scene Prediction (NSP)
$\mathbf{Z}_{v_{t}}=\texttt{SigLIP}(\mathbf{X}_{v_{t}}),\ \ \mathbf{H}_{v_{t}}=\texttt{MLP}(\mathbf{Z}_{v_{t}}),$ (1)
@@
Considering the critical role of both vehicle dynamics and high-level driving intent, we introduce a context encoder, a two-layer MLP module, to embed the ego vehicle’s current velocity $\boldsymbol{v}$, acceleration $\boldsymbol{a}$, and navigation command cmd into a context representation:
@@ 3.2 Decision Chain-of-Thought (DeCoT)
As shown in Fig. 2, the system prompt and navigation instructions $\mathbf{X}_{p}$ are first tokenized into token IDs and then encoded into textual embeddings $\mathbf{H}_{p}^{L_{p}\times D_{p}}$ using a text encoder, where $L_{p}$ and $D_{p}$ denote the number of text tokens and the dimensionality of the textual space, respectively. To enable effective multimodal alignment, we introduce a special <image> token into the tokenizer vocabulary. To preserve spatial awareness and allow the model to distinguish among different viewpoints, we explicitly annotate each token with its corresponding camera perspective in the prompt (e.g., CAM_FRONT: <image>, …, CAM_BACK: <image>). These <image> tokens are dynamically substituted with the corresponding visual embeddings processed by the image encoder, enabling seamless integration of textual and visual modalities. To better support the implementation of NSP and DeCoT, we introduce six additional special tokens beyond <image>. The input and target sequences for the LLM are structured as follows:
[…]
@@
Unlike traditional multi-turn QAs paradigms, ReasonPlan performs multi-step reasoning in a single forward pass, effectively leveraging the common sense reasoning capabilities of LLM for autonomous driving tasks.
@@ 4 PDR: Planning-oriented Decision Reasoning Dataset
To construct a reliable reasoning dataset, we extend the ground-truth annotations provided by Bench2Drive [26] with structured reasoning labels. Each reasoning trace has undergone thorough human verification to ensure consistency and interpretability across diverse scenarios. As shown in Fig. 4, the reasoning process comprises the following stages: Scene Understanding, Traffic Sign Recognition, Critical Object Identification for Risk Assessment, and Meta Action. In Appendix A, we detail the statistics of PDR and the annotation pipeline for each component.
@@ 5.1 Benchmark and Evaluation Metrics
Benchmark. We evaluate the closed-loop driving performance of ReasonPlan on Bench2Drive [26], which features challenging interactive scenarios based on the Carla leaderboard v2. To further assess its reasoning capabilities, we also conduct zero-shot evaluations on DriveOcclusionSim (DOS) [44], a suite of complex scenarios requiring models to infer global context from dynamic driving environments. For ablation studies, since evaluating 220 routes of Bench2Drive could take days, we use the Dev10 [26] for quick validation.
@@ 5.2 Comparison with State-of-the-arts
SOTA Performance in both Open-Loop and Closed-Loop Evaluations. As reported in Tab. 1, ReasonPlan achieves the best open-loop performance, reducing the L2 error to 0.61, demonstrating superior future trajectory prediction accuracy. Besides, in closed-loop evaluation, ReasonPlan delivers competitive performance compared to SOTA method DriveAdapter [46] which utilizes privileged expert feature distillation. Specifically, it achieves a DS of 64.01, significantly outperforming the non-distillation IL-based SOTA method, MomAD [23] by 16.1 (+33.6%). Furthermore, ReasonPlan improves the success rate by 16.44% (+90.78%) compared to MomAD [23]. ReasonPlan also achieves the highest efficiency score of 180.64 among all evaluated methods, demonstrating an effective and proactive driving policy. The comfort score of 25.63 reflects a common trade-off between trajectory agility and smoothness. Nonetheless, the comfort margin remains within acceptable bounds and could be further optimized via post-smoothing or low-level controller tuning. Furthermore, Tab. 1 presents the performance of ReasonPlan in the Multi-Ability evaluation across a diverse set of driving scenarios. ReasonPlan achieves a Mean Ability score of 36.66%, significantly outperforming all E2E baselines without leveraging any expert-feature distillation. These results underscore the model’s strong capability to reason across diverse driving intents and validate its robustness under complex, multi-intent scenarios.
@@ 6 Conclusion
In this work, we present ReasonPlan, a novel fine-tuning framework that adapts MLLMs for complex closed-loop scenarios. ReasonPlan introduces a temporally self-supervised next scene prediction task and explicit supervision over the decision chain-of-thought process, enabling unified integration of visual and textual modalities for interpretable planning. Through comprehensive evaluations on Bench2Drive and DOS, ReasonPlan achieves excellent performance in both open-loop and closed-loop settings. Notably, it demonstrates strong zero-shot generalization on unseen tasks, underscoring its robustness and potential for real-world deployment. Our results highlight the promise of MLLM-based frameworks in bridging high-level reasoning and low-level planning, paving the way for more cognitive and generalizable autonomous driving systems.

### v2

@@ Abstract
Due to the powerful vision-language reasoning and generalization abilities, multimodal large language models (MLLMs) have garnered significant attention in the field of end-to-end (E2E) autonomous driving. However, their application to closed-loop systems remains underexplored, and current MLLM-based methods have not shown clear superiority to mainstream E2E imitation learning approaches. In this work, we propose ReasonPlan, a novel MLLM fine-tuning framework designed for closed-loop driving through holistic reasoning with a self-supervised Next Scene Prediction task and supervised Decision Chain-of-Thought process. This dual mechanism encourages the model to align visual representations with actionable driving context, while promoting interpretable and causally grounded decision making. We curate a planning-oriented decision reasoning dataset, namely PDR, comprising 210k diverse and high-quality samples. Our method outperforms the mainstream E2E imitation learning method by a large margin of $16.44\%$ L2 and $16.1$ driving score on Bench2Drive benchmark. Furthermore, ReasonPlan demonstrates strong zero-shot generalization on unseen DOS benchmark, highlighting its adaptability in handling zero-shot corner cases. Code and dataset will be found in https://github.com/Liuxueyi/ReasonPlan.
@@ 1 Introduction
Recent efforts have explored dual-system frameworks [13, 14, 15], LLM distillation for enhancing E2E driving [16, 17], and direct trajectory prediction in textual form [18, 19, 20]. While promising, these approaches predominantly operate in open-loop settings or exhibit suboptimal performance in closed-loop evaluations. This limitation stems from their inability to perform context-aware reasoning and robust planning in closed-loop scenarios, where continuous adaptation to dynamic environments is essential [21]. We conclude three key challenges that limit the full exploitation of MLLMs’ reasoning capabilities: (1) Insufficient utilization of visual information. Existing methods [22] supervise the perception and decision-making process using only text or simply use predicted images as input for action generation [23], leading to constrained scene comprehension and the accumulation of compounded errors. (2) No explicit reasoning process. Prior works [18, 19] utilize multi-turn QAs fine-tuning LLMs to enhance the instruction-following ability, but they fall short of engaging the models’ Chain-of-Thought (CoT) reasoning capabilities. (3) Lack of planning-oriented high-quality reasoning datasets. Consequently, the full potential of MLLMs for E2E closed-loop planning remains underexplored, and current MLLM-based methods have yet to demonstrate clear superiority over imitation learning techniques [1, 2, 24] in closed-loop benchmarks.
@@
From the dataset perspective, we construct a large-scale instruction dataset tailored for closed-loop planning, called PDR, which contains 203,353 training samples and 11,047 testing samples. Using an automated annotation pipeline, PDR captures the complete decision reasoning process in training scenarios on the Bench2Drive [27], including the following stages: Scene Understanding, Traffic Sign Recognition, Critical Object Identification for Risk Assessment, and Meta Action. The dataset will be released publicly, serving as a foundation for learning structured and causally grounded decision reasoning.
ReasonPlan achieves a driving score of 64.01 and reduces the L2 error by 16.44% on Bench2Drive [27], outperforming prior state-of-the-art E2E imitation learning models. Furthermore, it demonstrates strong zero-shot generalization on DOS [28], highlighting its adaptability in decision-critical corner cases.
Our key contributions are as follows. (1) We propose ReasonPlan, a novel MLLM fine-tuning framework for complex closed-loop driving scenarios. The integration of NSP and DeCoT effectively couples the visual and language modalities, enabling a comprehensive decision reasoning process. (2) We construct PDR, a large-scale decision reasoning dataset via an automated annotation pipeline tailored for closed-loop planning, containing 210k diverse and high-quality samples. (3) As shown in Fig. 1, ReasonPlan demonstrates superior performance on the Bench2Drive under both open-loop and closed-loop settings, and shows strong zero-shot generalization on DOS scenarios.
@@ 3.1 Self-supervised Next Scene Prediction (NSP)
$\mathbf{Z}_{v_{t}}=\texttt{SigLIP}(\texttt{AnyRes}(\mathbf{X}_{v_{t}})),\ \ \mathbf{H}_{v_{t}}=\texttt{MLP}(\mathbf{Z}_{v_{t}}),$ (1)
@@
Considering the critical role of both vehicle dynamics and high-level driving intent, we introduce a context encoder, a two-layer MLP module, to embed the ego vehicle’s current velocity $\boldsymbol{v}$, acceleration $\boldsymbol{a}$, and driving command cmd (e.g. turn left) into a context representation:
@@ 3.2 Decision Chain-of-Thought (DeCoT)
As shown in Fig. 2, the system prompt and navigation instructions $\mathbf{X}_{p}$ are first tokenized into token IDs and then encoded into textual embeddings $\mathbf{H}_{p}^{L_{p}\times D_{p}}$ using a text encoder, where $L_{p}$ denotes the number of text tokens. To enable effective multimodal alignment, we introduce a special <image> token into the tokenizer vocabulary. To preserve spatial awareness and allow the model to distinguish among different viewpoints, we explicitly annotate each token with its corresponding camera perspective in the prompt (e.g., CAM_FRONT: <image>, …, CAM_BACK: <image>). These <image> tokens are dynamically substituted with the corresponding visual embeddings processed by the image encoder, enabling seamless integration of textual and visual modalities. To better support the implementation of NSP and DeCoT, we introduce six additional special tokens beyond <image>. The input and target sequences for the LLM are structured as follows:
$\texttt{User: \{velocity\} \{acceleration\} \{navigation command\} \{image tokens\}\_t \{prompt\}}$}.
@@
Unlike traditional multi-turn QAs paradigms, ReasonPlan performs multi-step reasoning in a single forward pass, effectively leveraging the commonsense reasoning capabilities of LLM.
@@ 4 PDR: Planning-oriented Decision Reasoning Dataset
To construct a reliable reasoning dataset, we extend the ground-truth annotations provided by Bench2Drive [27] with structured reasoning labels. The reasoning traces have undergone thorough human verification to ensure consistency and interpretability across diverse scenarios. As shown in Fig. 4, the reasoning process comprises the following stages: Scene Understanding, Traffic Sign Recognition, Critical Object Identification for Risk Assessment, and Meta Action. In Appendix A, we detail the statistics of PDR and the annotation pipeline for each component.
@@ 5.1 Benchmark and Evaluation Metrics
Benchmark. We evaluate the closed-loop driving performance of ReasonPlan on Bench2Drive [27], which features challenging interactive scenarios based on the Carla leaderboard v2. To further assess its reasoning capabilities, we also conduct zero-shot evaluations on DriveOcclusionSim (DOS) [28], a suite of complex scenarios requiring models to infer global context from dynamic driving environments. For ablation studies, we use the Dev10 [27] for quick validation.
@@ 5.2 Comparison with State-of-the-arts
SOTA Performance in both Open-Loop and Closed-Loop Evaluations. As reported in Tab. 1, ReasonPlan achieves the best open-loop performance with 0.61 L2 error, demonstrating superior trajectory prediction accuracy. In closed-loop evaluation, ReasonPlan delivers competitive performance compared to SOTA method DriveAdapter [47] which utilizes privileged expert feature distillation. Specifically, it achieves a DS of 64.01, significantly outperforming the non-distillation IL-based SOTA method, MomAD [24] by 16.1. Furthermore, ReasonPlan improves the success rate by 16.44% compared to MomAD [24]. ReasonPlan also achieves the highest efficiency score of 180.64 among all evaluated methods, demonstrating an effective and proactive driving policy. The comfort score of 25.63 reflects a common trade-off between trajectory agility and smoothness. Nonetheless, the comfort margin could be further optimized via post-smoothing or low-level controller tuning. Besides, ReasonPlan achieves a Mean Ability score of 36.66% across diverse driving scenarios, significantly outperforming all E2E baselines. These results underscore the model’s strong capability to reason across diverse driving intents and validate its robustness under complex, multi-intent scenarios.
@@ 6 Conclusion
In this work, we present ReasonPlan, a novel fine-tuning framework that adapts MLLMs for complex closed-loop scenarios. ReasonPlan introduces a self-supervised Next Scene Prediction task and explicit Decision Chain-of-Thought process, enabling unified integration of visual and textual modalities for interpretable planning. Through comprehensive evaluations on Bench2Drive and DOS, ReasonPlan achieves excellent performance in both open-loop and closed-loop settings. Notably, it demonstrates strong zero-shot generalization on unseen tasks, underscoring its robustness and potential for real-world deployment. Our results highlight the promise of MLLM-based frameworks in bridging high-level reasoning and low-level planning, paving the way for more cognitive and generalizable autonomous driving systems.
@@ 7 Limitations
8 Acknowledgement
This work was supported by the National Key Re-search and Development Program of China under Grant 2022YFA1004000, in part by Beijing Natural Science Foundation under Grant L253007 and 4242052 , in part by the State Key Laboratory of Safety Intelligent Mining in Non-coal Open-pit Mines, National Mine Safety Administration (Grant No.2024-ZD04).